Mr. Tanaka is 35 yrs old, and his son is 5. In how many yrs will MR. TANAKA be three times as old as his son?
Answer by ankor@dixie-net.com(22740)

. Tanaka is 35 yrs old, and his son is 5. In how many yrs will MR. TANAKA be three times as old as his son?
:
let y = no. of years for this to be true
y + 35 = 3(y+5)
y + 35 = 3y + 15
35 - 15 = 3y - y
20 = 2y
y = 20/2
y = 10 yrs
:
:
In 10 yrs T will be 45 and son will be 15, (3 times as old as his son)